A 36.21 g solid sample of mercury is heated from -39 oC its melting point, completely melted then heated to 92 oC, How much energy was absorbed by the sample? (Answer in kJ)

S=0.14 J/goC Delta Hf = 0.011 kJ/g


1
Expert's answer
2021-03-24T06:41:23-0400

Solution:

This problem can be summarized thusly:

1) A sample of solid mercury is melted at -39°C. The heat absorbed is calculated by multiplying the grams of solid mercury by the heat of fusion (ΔHf) in kJ/g.

q1 = ΔHf × m = 0.011 kJ/g × 36.21 g = 0.3983 kJ

q1 = 0.3983 kJ


2) A sample of liquid mercury is heated from -39°C to 92°C. The heat absorbed is calculated by using the specific heat of liquid mercury (S = 0.14 J/g°C) and the equation: q = m × S × ΔT.

q2 = m × S × (Tf - Ti)

q2 = 36.21 g × 0.14 J/g°C × (92 - (-39))°C = 664.09 J = 0.6641 kJ

q2 = 0.6641 kJ


q = q1 + q2 = 0.3983 kJ + 0.6641 kJ = 1.0624 kJ = 1.06 kJ

q = 1.06 kJ


Answer: 1.06 kJ of energy was absorbed by the sample.